OpenStudy (ahrimain):

18(3+14) x 6(21+4) Without Distribution

OpenStudy (pawanyadav):

3+14=17 And 21+4=25 So 18(3+14) x 6(21+4) =18(17) × 6(25) Can you multiply it now?